OpenStudy (anonymous):

How would you simplify (4x)/(9)+(x-10)/(9) showing work?

OpenStudy (anonymous):

5x-10/9

OpenStudy (anonymous):

showing work? or its just a straight answer?

OpenStudy (anonymous):

they have a comon denomonator so add 4x+x-10/9 then simplify

OpenStudy (anonymous):

can you show it so i can understand how to do the next problems as well

OpenStudy (anonymous):

i just did kay 4x/9+x-10/9 <<<since they have a common denominator just add the numerator 4x+x-10 if they both aren't over the same number then find a comon denominator

OpenStudy (lalaly):

\[\frac{4x}{9} + \frac{(x-10)}{9} = \frac{4x+x-10}{9} = \frac{5x-10}{9}\]

OpenStudy (anonymous):

^^^^ haha thanks i wasn't sure how to do that so it wasv confusing, but thats what i was saying

OpenStudy (anonymous):

Thanks guys!!!!